OpenStudy (anonymous):

MO bisects angle LMN, m ∠ LMN = 6x - 28, m ∠ LMO = x + 34. Find m ∠ NMO.

OpenStudy (anonymous):

58 right?

OpenStudy (raden):

i cant login, can you make a diagram here ?

OpenStudy (anonymous):

Can you help me with another? I figured this one out.

OpenStudy (raden):

i'll try

OpenStudy (anonymous):

If T is the midpoint of line SU, what are ST, TU, and SU? ST = 20, TU = 20, and SU = 40 ST = 80, TU = 80, and SU = 160 ST = 16, TU = 80, and SU = 160 ST = 102, TU = 102, and SU = 204

OpenStudy (anonymous):



OpenStudy (raden):

oke, because T is the midpoint of line SU, then satisfies : ST = TU 5x = 3x + 32 (solve for x) 5x - 3x = 32 2x = 32 x = 32/2 x = 16 now subtitute the value of x = 16 into the equations of ST = 5x = 5 (16) = 80 and TU = 3x + 32 = 3(16) + 32 = 48 + 32 = 80 also SU = ST + TU = 80 + 80 = 160

OpenStudy (anonymous):

So 16, 80, 160?

OpenStudy (raden):

no, it is ST = 80, TU = 80, and SU = 160

OpenStudy (anonymous):

Ohh! I understand now because 80 + 80..